OpenStudy (aravindg):

time is of a particle is given by x^2 +x (x>0) find its retardation

OpenStudy (jamesj):

retardation = negative of acceleration?

OpenStudy (aravindg):

ya

OpenStudy (jamesj):

so if s(x) = x^2 + x then velocity v(t) = ds/dt and acceleration a(t) = d^2s/dt^2 So find the second derivative.

OpenStudy (jamesj):

Wait. The position of a particle at time t is s(t) = t^2 + t. Or is the question saying something else?

OpenStudy (aravindg):

time

OpenStudy (jamesj):

So you're saying that x is the particle's displacement at it satisfies the relation that at time t: t = x^2 + x ?

OpenStudy (aravindg):

yes

OpenStudy (jamesj):

Ok. In that case, solve this equation for x and then differentiate it. So you tell me: what is x as a function of t?

OpenStudy (aravindg):

quadratic roots?

OpenStudy (jamesj):

x^2 + x - t = 0. Solve for x

OpenStudy (jamesj):

Found the expression for x yet?

OpenStudy (jamesj):

Quadratic formula. a = 1, b = 1, c = -t